If a child on a bicycle can accelerate to an average rate of 0.53 m/s(squared), how long would it take to increase the bicycles velocity from 0.68m/s [N] to 0.89m/s [N]
1
Expert's answer
2018-10-18T11:10:09-0400

We can find the time that the child takes to increase the bicycle’s velocity from the kinematic equation:

v=v_0+at,

t=(v-v_0)/a=(0.89 m/s-0.68 m/s)/(0.53 m/s^2 )=0.39 s.